OpenStudy (anonymous):

how do you solve this x-1=x+1 --- ---- 2x+1 x-1

OpenStudy (radar):

I think you end up with (x+1)/(2x+1) providing the equation is what I think it is

OpenStudy (radar):

I am assuming it is: (x-1)/(2x+1) = (x+1)/(x-1)

OpenStudy (anonymous):

yes

OpenStudy (radar):

O.K. The (x-1) can be eliminated.

OpenStudy (radar):

They "cancel" but I can't remember the rule. I'm thinking.

OpenStudy (anonymous):

The (x-1)'s can't be eliminated if you have an = sign between the fractions. If it is a multiplication problem, then they could be eliminated.

OpenStudy (radar):

ooops

OpenStudy (anonymous):

You need to cross multiply. (x - 1)(x -1) = (2x + 1)(x + 1)

OpenStudy (anonymous):

x^2 - 2x + 1 = 2x^2 + 3x + 1

OpenStudy (anonymous):

make this equal to 0 0 = x^2 + 5x

OpenStudy (anonymous):

Factor 0 = x(x + 5) So x = 0 or x = -5

OpenStudy (anonymous):

thanks i can see exactly were you got every number from ill do the next one and see if i understood it

OpenStudy (radar):

I am trying to prove those answers, when I subsstitute the -5 it does not check out. Where am I going wrong?

OpenStudy (radar):

(-5-1)/(2(-5)+1)=(-5+1)/(-5-1) -6/-9 = -4/-6 Unh unh Something is wrong with the solution.

OpenStudy (radar):

Never mind I didn't reduce!! lol

OpenStudy (radar):

2/3=2/3

OpenStudy (radar):

I had a senior moment!